OpenStudy (anonymous):

Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y

OpenStudy (anonymous):

A. no B. yes; k = -5/3 C. yes; k = 5/3 D. yes; k = 3/5 choices

OpenStudy (anonymous):

Yes, can write as -3y = 5x y = (-5/3) *x the constant of variation is -5/3

OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2?

OpenStudy (anonymous):

A. y = –0.75x; 1.50 B. y = –1.33x; 2.67 C. y = 1.33x; –2.67 D. y = 0.13x; –0.25 choices

OpenStudy (anonymous):

so the answer is B B. y = –1.33x; 2.67

OpenStudy (anonymous):

Y=bx when x = -6 y=b*-6 but we are given that when b =-6, y =8 Therfore rewrite: y=B*(-6) as 8 = b*(-6) => b = -8/6 Therfore the equation is y= (-8/6)*x = -1.333x y= -1.333x b) when x=-2 just substitute x =-2 in the above equation: y= -1.333x , plug in -2 for x y= -1.33 * (-2) = 2.666 ( multiplying two negative signs turns to a positive sign) => y= 2.666 rounded to 2.67 so the answer is B B. y = –1.33x; 2.67